OpenStudy (anonymous):

I'm very confused, can someone help explain? Your schools talent show will feature 12 solo acts and ensemble 2 acts. The show will last 90 min. The 6 solo performers judged best will give a repeat performance at a second 60-min show, which will also feature the 2 ensemble acts. Each solo lasts x minutes, and each ensemble act lasts y minutes. A. Write a system of equations to model the situation. B. Solve the system from part(a). How long is each solo act? How long is each ensemble act?

OpenStudy (anonymous):

12x+2y=90

OpenStudy (anonymous):

6x+2y=60

OpenStudy (anonymous):

second eq. 2y=60-6x 12x+(60-6x)=90 12x-6x=90-60 6x=30 x=30/5=5 min 2y=60-6(5) 2y=60-30 2y=30 y=30/2=15 min

OpenStudy (funnyguy2):

^^^^

OpenStudy (anonymous):

thanks!